OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4.

OpenStudy (souvik):

try to get this equation like this ...\[\huge{y=\color\red{m}x+\color\green{b}}\] here m is the slope and b is the y intercept....

OpenStudy (anonymous):

slope = -1; y intercept at (0, -negative four thirds) slope = 1; y intercept at (0, -negative four thirds) slope = 3; y intercept at (0, -1) slope = -3; y intercept at (0, 1)

OpenStudy (texaschic101):

do you understand what souvik just explained ? You can find your answer by following what he said :)

OpenStudy (anonymous):

yeah I get it

OpenStudy (texaschic101):

3x - 3y = -4 get your 3y on one side and everything else on the other side

OpenStudy (texaschic101):

do you want to work it out together ?

OpenStudy (texaschic101):

guess not.....I tried

OpenStudy (anonymous):

Yes I do sorry

OpenStudy (anonymous):

y=3x−4 all over 3

OpenStudy (anonymous):



OpenStudy (texaschic101):

3x - 3y = -4 -3y = -3x - 4 (multiply by -1 to make -3y positive) 3y = 3x + 4 (now divide by 3) y = 3/3x + 4/3 y = x + 4/3

OpenStudy (anonymous):

the answer is -1 and -4/3 right

OpenStudy (texaschic101):

no

OpenStudy (souvik):

the answer is 1 and (0,4/3) as m=1 and b=4/3

OpenStudy (texaschic101):

look at how I worked out the problem.....up above

OpenStudy (texaschic101):

no...it is slope 1 and y intercept 4/3

OpenStudy (anonymous):

that's not an option

OpenStudy (anonymous):

slope = -1; y intercept at (0, -negative four thirds) slope = 1; y intercept at (0, -negative four thirds) slope = 3; y intercept at (0, -1) slope = -3; y intercept at (0, 1) these are the options

OpenStudy (texaschic101):

your options are incorrect

OpenStudy (anonymous):

tell that to flvs.... I cant do anything about it

OpenStudy (texaschic101):

I don't know what answer you should put because the answers are all incorrect

OpenStudy (texaschic101):

did you write the problem correctly

OpenStudy (souvik):

@texaschic101 the equation is 3x-3y=4 but u took it 3x-3y=-4